# Cross-Modality Perturbation Synergy Attack for Person Re-identification

### 18 Jan 2024 v1

@@ 4 Experiments
AGW baseline TPAMI 47.50 84.39 92.14 47.65 54.17 91.14 95.98 62.97
M-FGSM attack TPAMI 25.79 49.04 57.96 19.24 20.56 38.91 46.35 15.89
ODFA attack IJCV 25.43 47.49 56.38 19.00 14.62 29.92 36.42 11.35
ours attack —— 1.11 8.67 16.14 1.41 1.31 7.47 10.36 1.23
DDAG baseline ECCV 54.75 90.39 95.81 53.02 61.02 94.06 98.41 67.98
M-FGSM attack TPAMI 28.36 52.47 60.76 23.11 24.85 40.74 49.22 18.40
ODFA attack IJCV 27.75 50.26 59.14 22.30 17.62 32.64 40.03 14.83
ours attack —— 1.62 7.59 14.46 1.84 1.45 7.71 10.72 1.25
Table 2: Results for attacking cross-modality ReID systems on the RegDB dataset. It reports on visible images querying infrared images and vice versa. Rank at $r$ accuracy (%) and mAP (%) are reported.
@@
AGW baseline TPAMI 70.05 86.21 91.55 66.37 70.49 87.21 91.84 65.90
ours attack —— 2.29 9.06 18.35 3.92 1.93 11.44 19.30 3.46
DDAG baseline ECCV 69.34 86.19 91.49 63.46 68.06 85.15 90.31 61.80
ours attack —— 1.33 10.28 19.06 3.79 1.35 9.52 17.52 3.19

### 19 Jan 2024 v2

@@ 4 Experiments
AGW baseline TPAMI 2022 47.50 84.39 92.14 47.65 54.17 91.14 95.98 62.97
M-FGSM attack TPAMI 2020 25.79 49.04 57.96 19.24 20.56 38.91 46.35 15.89
ODFA attack IJCV 2023 25.43 47.49 56.38 19.00 14.62 29.92 36.42 11.35
ours attack ———— 1.11 8.67 16.14 1.41 1.31 7.47 10.36 1.23
DDAG baseline ECCV 2020 54.75 90.39 95.81 53.02 61.02 94.06 98.41 67.98
M-FGSM attack TPAMI 2020 28.36 52.47 60.76 23.11 24.85 40.74 49.22 18.40
ODFA attack IJCV 2023 27.75 50.26 59.14 22.30 17.62 32.64 40.03 14.83
ours attack ———— 1.62 7.59 14.46 1.84 1.45 7.71 10.72 1.25
Table 2: Results for attacking cross-modality ReID systems on the RegDB dataset. It reports on visible images querying thermal images and vice versa. Rank at $r$ accuracy (%) and mAP (%) are reported.
@@
AGW baseline TPAMI 2022 70.05 86.21 91.55 66.37 70.49 87.21 91.84 65.90
ours attack ———— 2.29 9.06 18.35 3.92 1.93 11.44 19.30 3.46
DDAG baseline ECCV 2020 69.34 86.19 91.49 63.46 68.06 85.15 90.31 61.80
ours attack ———— 1.33 10.28 19.06 3.79 1.35 9.52 17.52 3.19